### v1

@@ Abstract
We analyze the convergence properties of gradient descent algorithms on Riemannian manifolds. We study randomization of the tangent space directions of Riemannian gradient flows for minimizing smooth cost functions (of Morse–Bott type) to obtain convergence to local optima. We prove that through randomly projecting Riemannian gradients according to the Haar measure, convergence to local optima can be obtained almost surely despite the existence of saddle points. As an application we consider ground state preparation through quantum optimization over the unitary group. In this setting one can efficiently approximate the Haar-random projections by implementing unitary 2-designs on quantum computers. We prove that the respective algorithm almost surely converges to the global minimum that corresponds to the ground state of a desired Hamiltonian. Finally, we discuss the time required by the algorithm to pass a saddle point in a simple two-dimensional setting.
@@ 1 Introduction
Both gradient systems and numerical linear algebra come with a long history as well as a vibrant research activity. For instance, gradient systems have been advanced lately from Hilbert manifolds [47] to general metric spaces (e.g., in the context of optimal mass transport [2, 45]). In contrast, numerical linear algebra has faced factorization problems of huge tensor structures [6] (tensor SVD) using “randomization strategies” for their classical algorithms in order to handle these amounts of data. Only a few decades ago a systematic study of their interplay started by analyzing the QR-algorithm and interior point methods from a (Riemannian) geometric point of view, cf. the collection of [11]. In this development the seminal work of [12, 13] on the so-called “double-bracket flow” on orbits of the orthogonal group (and likewise of the unitary group in [14]), followed by independent similar ideas of [16], has sparked a plethora of applications. For an early overview on optimization techniques on Riemannian manifolds (including higher-order methods) see [56] in [11]. A mathematical account in all detail can be found in the monograph by [28] including discretization schemes, suitable step sizes, and proofs of convergence. For the latter, the authors heavily exploited the Morse-Bott structure of the cost function (i.e. the particular stucture of the Hessian at possibly non-isolated critical points, cf. Subsec. 2.3 and [22] to obtain convergence to a single critical point. Based on these developments, higher-order methods with applications to further cost functions on classical matrix manifolds are treated in [1]. In [54] these ideas have been extended to flows and their discretization schemes on more general homogeneous spaces. There they were also connected to applications in quantum dynamics as well as to applications in numerical (multi-)linear algebra like higher-rank tensor approximations—as elaborated by [19] to complement standard methods (like higher-order powers of [34] or quasi Newton methods of [53]).
@@
Often gradient-flow approaches for deriving optimization schemes on Riemannian manifolds hinge on the computability of the Riemannian exponential ensuring to take the gradient from the tangent space back on the manifold. To be precise, this is crucial whenever the manifold cannot be identified with its tangent spaces such as for the unitary group or unitary orbits.
@@
With these stipulations, Riemannian optimization techniques constitute a versatile toolbox, since a cost function can readily be tailored to the particular application such as finding extremal eigen- or singular values (i.e. problems otherwise addressed by variational approaches). These instances also connect to other branches of mathematics such as numerical and $C$-numerical ranges and their extremal values, the numerical and $C$-numerical radii (see [27, 38]).
Over the last decade, techniques from Riemannian optimization have also been adopted in quantum information science. For example, in quantum computing, hybrid quantum-classical algorithms [8], such as variational quantum algorithms (VQAs) [15, 59], have been developed to solve ground state problems [49] (i.e. typically minimal eigenvalue problems), combinatorial optimization [23], and quantum machine learning problems [9]. In VQAs a fixed parameterized quantum circuit, i.e. a parameterized set of unitary transformations (usually implemented by a fixed set of unitary gates), is iteratively optimized in tandem with a classical computer to minimize a cost function whose global optima encode the solution(s) to the desired problem. More precisely, a quantum processor, including a measurement device, is used to output the current (e.g., expectation) value of the cost function in each iterative step. Interleaved, the classical computer component provides the optimization routine that determines the parameter update for the next iteration on the quantum processor. Since quantum computers can, in principle, handle some classically exponentially complex problems, the hope is that VQAs will be able to solve problems of large size a solely classical algorithm would be incapable of.
However, since the optimization problems VQAs aim to solve are typically non-convex, which can often be traced back to the quantum circuit parameterization [41, 36], the iterative search for the optimal parameters can get stuck in suboptimal solutions [10].
To overcome this challenge, adaptive quantum algorithms have been designed. Instead of fixing a quantum circuit and picking a parameterization as in VQAs, adaptive algorithms successively grow a quantum circuit informed by measurement data from the quantum processor. In this approach the cost function is minimized while the quantum circuit is grown, see [25, 63, 42, 43]. One of the most prominent examples of such a strategy is the so-called ADAPT-VQE algorithm [25], which was originally proposed to solve ground state problems in chemistry, and was further developed later to solve combinatorial optimization problems on quantum computers [58].
Unfortunately, since typically in adaptive quantum algorithms the circuit growth is informed by gradient estimates performed by the quantum computer, adaptive quantum algorithms face similar challenges as VQAs, i.e., the adaptive circuit growth can get stuck when gradients vanish. This issue has recently been addressed by identifying adaptive quantum algorithms as quantum computer implementations of Riemannian gradient flows on the special unitary group $\text{SU}(d)$ of dimension $d=2^{n}$ where $n$ is the number of qubits of the quantum computer [63]. Since implementing Riemannian gradient methods on quantum computers requires in general exponential resources (i.e., the number of gates and/or the number of iterations of the corresponding quantum circuit growths exponentially in $n$). [63] proposed projecting the Riemannian gradient into smaller dimensional subspaces that scale polynomially in $n$, which in turn yields scalable quantum computer implementations. However, the proposed “dimension reduction” strategy of the Riemannian gradient to obtain efficient quantum computer implementations comes with a cost: While it can be shown that in the quantum computing setting (i.e., the Riemannian gradient flow on the special unitary group for grounds state problems) the full gradient flow converges almost surely to the problem solution given by the global minimizer, despite the existence of strict saddle points (i.e., saddle points at which the Hessian has at least one negative eigenvalue) [37], for the projected versions no convergence guarantees can be made. Indeed, numerical simulations suggest that existing versions of gradient algorithms that are efficiently implementable on quantum computers also suffer from converging to suboptimal solutions. Inspired by the success of classical randomized gradient descent [52], this problem has most recently been addressed by some of the authors in [43] via “random projections” of the Riemannian gradient, as depicted in Fig. 1, which are efficiently implementable on quantum computers. An efficient quantum computer implementation of these random directions can be achieved through exact and approximate unitary 2-designs [21]. The numerical experiments presented in [43] indicate that such a randomization procedure of the Riemannian gradient indeed yields (almost surely) convergence to the problem solution.
@@
In this paper we consider the randomized gradient algorithm presented in [43] in the more general setting of Morse–Bott cost functions on Riemannian manifolds. For both continuous and discrete probability distributions of the random gradient directions we prove that the algorithm almost surely converges to a local minimum even. Finally we show how the result can be applied to quantum optimization tasks, such as ground state preparation. In the quantum setting there are no local optima, thereby we prove that the randomized gradient algorithm converges almost surely to the global minimum.
@@
Figure 1: Schematic representation of the randomized gradient descent algorithm on a Riemannian manifold $\mathcal{M}$. At a point $X_{k}$ on $\mathcal{M}$ the Riemannian gradient $\operatorname{grad}J(X_{k})$ of a cost function $J:\mathcal{M}\to\mathbb{R}$ lies in the tangent space $T_{X_{k}}\mathcal{M}$ depicted as plane spanned by $\Omega_{k}$ and its orthocomplement $\Omega_{k}^{\perp}$. Instead of iteratively following the full gradient in order to optimize $J$, here in each step we consider a projection of $\text{grad}J(X_{k})$ onto a randomly chosen tangent-space direction $\Omega_{k}$. This is key for efficient implementations of gradient flows in large-scale quantum optimization problems. In the case of the special unitary group the randomized directions are given by (traceless) skew-Hermitian operators $\Omega_{k}=iH_{k}$. The corresponding quantum circuit consists of a sequence of unitary transformations as shown in Fig. 2.
Figure 2: Schematic implementation of the discretized Riemannian gradient scheme by a quantum algorithm in case the manifold is $\mathcal{M}=SU(2^{n})$ or $\mathcal{M}=Ad_{SU(2^{n})}$. For both instances the dimension of the problem scales exponentially with the number $n$ of quantum particles (here qubits). Each quantum circuit $U_{k}$ is adaptively turned into $U_{k+1}=\exp\left(\frac{-i\eta\,dJ(0)}{d\theta_{k}}H_{k}\right)U_{k}$ by moving in a random tangent-space direction $iH_{k}$ whose corresponding projection $\frac{dJ_{k}(0)}{d\theta_{k}}=\langle\text{grad}J(U_{k}),iH_{k}U_{k}\rangle$ is measured on a quantum device, where $\eta$ denotes the learning rate and we define the cost-function value at the $k$-th iteration as $J_{k}(x)\equiv J(x,U_{k})$ where $J(x,U_{k})=\text{tr}\{A\exp(-ixH_{k})U_{k}\rho U_{k}^{*}\exp(ixH_{k})\}$ (see Sec. 5). The measurement projection is done on the quantum computer, and the statistics of repeated measurements for the expectation value is done on the classical computer. In this way a randomized gradient flow as in Fig. 1 is implemented. In this work we show that for sufficiently small $\eta$ such a (discretized) flow converges almost surely to the global optimum.
@@
In Sec. 2 we introduce the fundamental notions from Riemannian geometry to describe gradient flows (Sec. 2.1), recollecting some well-known results on the convergence of gradient flows (Sec. 2.2), and their algorithmic implementation as gradient descents (Sec. 2.4). In Sec. 3 we describe a gradient descent algorithm where in each step the full gradient of a high-dimensional problem is projected on just a randomly chosen direction of the tangent space, considering both continuous and discrete distributions.
The main main result is obtained in Secs. 3.2 and 3.3, where we prove that a randomly projected gradient descent algorithm to a smooth Morse–Bott function (with compact sublevel sets) on a Riemannian manifold comes with two vitally beneficial properties: (i) it almost surely escapes saddle points—as is shown in Lem. 3.11 and as a consequence (ii) it almost surely converges a local minimum—as shown in Thm. 3.14.
In Sec. 4 the case of a two-dimensional saddle point is studied approximately using analytical methods and also using numerical simulation. We obtain good approximations for the time necessary to pass the saddle point.
Finally Sec. 5 shows how the previous results can be applied in quantum optimization to the problem of ground state preparation.
@@ 2.1 Basic Concepts
First we recall some basic notions and notations from Riemannian geometry. Let $M$ denote a finite dimensional smooth manifold of dimension $n$ with tangent and cotangent bundles ${\sf T}M$ and ${\sf T}^{*}M$, respectively. Moreover, let $M$ be endowed with a Riemannian metric $g$, i.e. a smoothly varying scalar product $g(\mathbf{x})=\langle\,\cdot,\cdot\,\rangle_{\mathbf{x}}$ on each tangent space ${\sf T}_{\mathbf{x}}M$, $\mathbf{x}\in M$. The pair $(M,g)$ is called a Riemannian manifold.
Every Riemannian manifold is equipped with a Riemannian density $\mu$ induced by the metric $g$, see [35, Prop. 2.44]. This density can be used to integrate functions, to induce a measure on $M$ and to define sets of measure zero. Equivalently (but without introduction a Riemannian density), one could say that $S\subset M$ has measure zero if there is an atlas such that $S\cap{\rm dom}\,\sigma$ has Lebesgue measure zero in every chart $\sigma$. In particular any submanifold of dimension strictly smaller than $N$ has measure zero.
@@ 2.2 Gradient Flows and Asymptotic Behaviour (I)
$df(\mathbf{x})\xi=\langle\operatorname{grad}f(\mathbf{x}),\xi\rangle_{\mathbf{x}}$ (1)
@@
In local coordinates the above reads as follows: Let $\sigma:U\to\mathbb{R}^{n}$ be a chart about $\mathbf{x}\in M$, denote $\mathbf{y}=\sigma(\mathbf{x})$, and let $\tilde{f}:=f\circ\sigma^{-1}$ be the chart representation of $f$. Moreover let $\tilde{g}:=(\sigma^{-1})^{*}g$ be the chart representation of the metric $g$ on $\sigma(U)$, i.e. $\tilde{g}(\mathbf{y})=\sum_{i,j=1}^{n}\tilde{g}_{ij}(\mathbf{y})\,dy^{i}dy^{j}$, where $\tilde{G}(y):=\big{(}\tilde{g}_{ij}(\mathbf{y})\big{)}$ is a positive definite matrix varying smoothly in $\mathbf{y}\in\sigma(U)$. Finally let $\tilde{F}:=(\sigma^{-1})^{*}F$ be the chart representation of the gradient vector field $\operatorname{grad}f$. Then it holds that
$\tilde{F}^{j}(\mathbf{y})=\sum_{i=1}^{n}\tilde{g}^{ij}(\mathbf{y})\frac{\partial\tilde{f}}{\partial y_{i}}(\mathbf{y}),$
where $\tilde{g}^{ij}(\mathbf{y})$ are the entries of the inverse of $\tilde{G}(\mathbf{y})$ and $\frac{\partial}{\partial y_{i}}$ denotes the $i$-th partial derivative.
Any convergence analysis of gradient systems is based on the following two observations: (i) the critical points of $f$ are the equilibria of (2), (ii) $f$ constitutes a Lyapunov function for (2), i.e. $f$ is monotonically decreasing along solutions. Despite their simplicity they immediately allow several non-trivial statements about the asymptotic behavior of (2) which can be characterized by its $\omega$-limit sets
@@
If f has compact sublevel sets, i.e. if the sets $\{\mathbf{x}\in M:f(\mathbf{x})\leq c\}$ are compact for all $c\in\mathbb{R}$, then every solution of (2) exists for $t\geq 0$ and its $\omega$-limit set is a non-empty, compact and connected subset of the set of critical points of $f$.
Although, Proposition 2.1 shows that $\omega(\mathbf{x})$ is contained in the set of critical points of $f$ it does not guarantee convergence to a single critical point and, indeed, there are smooth gradient systems the trajectories of which exhibit a non-trivial convergence behavior to the set of critical points (cf. “Mexican hat counter-example” by [18]). Yet for isolated critical points one has the following trivial consequence of Proposition 2.1.
@@
If f has compact sublevel sets and all critical points are isolated, then any solution of (2) converges (for $t\to\infty$) to a single critical point of $f$. Moreover if additionally all saddle points are strict111see Subsec. 2.3 below then for almost all initial points (in the sense of Sec. 2.1) the flow converges to a local minimum.

### v2

@@ Abstract
We analyze convergence of gradient-descent methods on Riemannian manifolds. In particular, we study randomization of Riemannian gradient algorithms for minimizing smooth cost functions (of Morse–Bott type). We prove that randomized gradient descent methods, where the Riemannian gradient is replaced by a random projection of it, converge to a single local optimum almost surely despite the existence of saddle points. We consider both uniformly distributed and discrete random projections. We also discuss the time required to pass a saddle point. As a major application, we consider ground-state preparation through quantum optimization over the unitary group. In mathematical terms our randomized algorithm applied to the trace function $U\to\operatorname{tr}(AU\rho U^{*})$ almost surely converges to its global minimum. The minimum corresponds to the smallest eigenvalue (ground state) of the selfadjoint operator $A$ (Hamiltonian) if $\rho$ is a rank-one projector (pure state). In this setting, one can efficiently replace the uniform random projections by implementing so-called discrete unitary 2-designs.
@@ 1 Introduction
Both gradient systems and numerical linear algebra come with a long history, as well as a vibrant research activity. For instance, gradient systems have been advanced lately from Hilbert manifolds [56] to general metric spaces (e.g., in the context of optimal mass transport [2, 53]). In contrast, numerical linear algebra has faced factorization problems of large tensor structures [8] (tensor SVD) using “randomization strategies” for their classical algorithms in order to handle these huge amounts of data. A few decades ago, a systematic study of their interplay started by analyzing the QR-algorithm and interior point methods from a (Riemannian) geometric point of view, cf. the collection of [13]. In this development, the seminal work of [15, 16] on the so-called “double-bracket flow” on orbits of the orthogonal group (and likewise of the unitary group in [17]), followed by independent similar ideas of [19], has sparked a plethora of applications. For an early overview on optimization techniques on Riemannian manifolds (including higher-order methods) see [65] and [13]. A detailed mathematical account can be found in the monograph by [33], including discretization schemes, suitable step sizes, and proofs of convergence. For the latter, the authors heavily exploited the Morse–Bott structure of the respective cost functions (i.e. the particular structure of the Hessian at possibly non-isolated critical points, cf. Subsec. 2.3 here and [26]) to obtain convergence to a single critical point. Based on these developments, higher-order methods with applications to a variety of cost functions on classical matrix manifolds are treated in [1] or, more recently, by [14]. In [63], the above ideas have been extended to flows and their discretization schemes on more general homogeneous spaces. There, the authors also established connections to applications in quantum dynamics as well as to applications in numerical (multi-)linear algebra, such as higher-rank tensor approximations — as elaborated by [23] to complement standard methods (such as higher-order powers of [40] or quasi-Newton methods of [62]).
@@
Often, gradient-flow approaches for deriving optimization schemes on Riemannian manifolds hinge on the computability of the Riemannian exponential, ensuring to take the gradient from the tangent space back to the manifold. To be precise, this is crucial whenever the manifold cannot be identified with its tangent spaces, such as for the unitary group and its orbits. But, in particular in this case, analytical expressions and efficient numerical approximations for the Riemannian exponential are well known.
@@
With these stipulations, Riemannian optimization constitutes a versatile toolbox, since cost functions can readily be tailored to the particular application such as finding extremal eigen- or singular values. These instances also connect to other branches of mathematics such as numerical and $C$-numerical ranges and their extremal values, the numerical and $C$-numerical radii (see [31, 45]).
Over the last decade, techniques from Riemannian optimization have also been adopted in quantum information science. For example, in quantum computing, hybrid quantum-classical algorithms [10], such as variational quantum algorithms (VQAs) [18, 69], have been developed to solve ground-state problems [58], combinatorial optimization problems [27], and quantum machine learning problems [11]. [73] proposed projecting the Riemannian gradient into smaller dimensional subspaces to obtain scalable [^1] Sec. 5 gives a precise meaning of ‘scalable quantum computers’ quantum-computer implementations. However, the “dimension reduction” strategy comes at a cost: While the full gradient flow converges almost surely to some local minimum in spite of strict saddle points222i.e., saddle points at which the Hessian has at least one negative eigenvalue [44], for the projected versions no such guarantees can be made. Indeed, numerical simulations see convergence to artificial local minima induced by the projection into a fixed subspace. Inspired by the success of classical randomized gradient descent [61], the problem of artificial minima has recently been circumvented by some of the authors in [50] via “random projections” of the Riemannian gradient. (Details on different optimization algorithms and their efficient implementations on quantum computers are postponed to Section 5.)
@@
In this paper we consider the randomized gradient algorithm presented in [50] in the more general setting of the large class [^3] Indeed, on a compact manifold, the set of Morse functions, which is a subset of the Morse–Bott functions, forms an open dense subset of the space of smooth functions [3]. In this sense, Morse–Bott functions are generic. of Morse–Bott cost functions on Riemannian manifolds. For both continuous and discrete probability distributions of the random gradient directions, we prove that in fact the algorithm almost surely converges to a local minimum. This is a significant improvement on the state-of-the-art, where thus far (see [32]) only convergence to the set of critical points has been proven without the Morse–Bott assumption. Related papers have established convergence to local optima for perturbed and stochastic gradient descent [36], also generalized to Riemannian manifolds [21] and [66]. Finally, we show how our result can be applied to quantum optimization tasks, such as ground state preparation: It is by construction that in the specific quantum setting, there are no local optima, whereby we prove that our randomized gradient algorithm then converges almost surely to a single global minimum.
@@
Figure 1: Schematic representation of a randomized gradient descent algorithm on a Riemannian manifold $M$. At a point $\mathbf{x}_{i}$ on $M$, the Riemannian gradient $\operatorname{grad}f(\mathbf{x}_{i})$ of a cost function $f:M\to\mathbb{R}$ lies in the tangent space ${\sf T}_{\mathbf{x}_{i}}M$ depicted as the plane spanned by $\mathbf{u}_{i}$ and its orthocomplement. Instead of iteratively following the full gradient in order to optimize $f$, here in each step we consider a projection of $\text{grad}f(\mathbf{x}_{i})$ onto a randomly chosen tangent-space direction $\mathbf{u}_{i}$. This is key for efficient implementations of gradient flows in large-scale quantum optimization problems. In the case of the special unitary group the randomized directions are given by (traceless) skew-Hermitian operators $\mathbf{u}_{i}=iH_{i}$. The corresponding quantum circuit consists of a sequence of unitary transformations as shown in Fig. 5, where the cost function is then denoted by $J$.
@@
In Sec. 2 we introduce the fundamental notions from Riemannian geometry to describe gradient flows (Sec. 2.1), recollecting some well-known results on the convergence of such gradient flows (Secs. 2.2 and 2.3), and their algorithmic implementation as gradient descents (Sec. 2.4). In Sec. 3 we describe a gradient descent algorithm where, in each step, the full gradient of a high-dimensional problem is projected on just a randomly chosen direction of the tangent space as illustrated in Fig.1. We consider both uniformly distributed444with respect to the Haar measure and discrete random directions.
The main results are obtained in Secs. 3.2 and 3.3, where we prove that a randomly projected gradient descent algorithm of a smooth Morse–Bott cost function (with compact sublevel sets) on a Riemannian manifold comes with two vitally beneficial properties: (i) it almost surely escapes saddle points — as is shown in Lem. 3.14, and as a consequence (ii) it almost surely converges to a single local minimum — as shown in Thm. 3.17.
In Sec. 4, the case of a two-dimensional saddle point is studied using analytical methods and numerical simulation. We obtain good approximations for the time necessary to pass the saddle point.
Finally, Sec. 5 shows how the previous results can be applied in quantum optimization: by construction, the important problem class of ground-state preparation comes with the particular advantage that our algorithm achieves almost sure convergence to a single global minimum. Moreover, in these cases the algorithm can be efficiently implemented on quantum computers.
@@ 2.1 Basic Concepts
[…]
Every Riemannian manifold is equipped with a Riemannian density $\mu$ induced by the metric $\mathsf{g}$, see [42, Prop. 2.44]. This density can be used to integrate functions, to induce a measure on $M$ and to define sets of measure zero. Equivalently (but without introduction of a Riemannian density), one could say that $S\subset M$ has measure zero if there is an atlas such that $S\cap{\rm dom}\,\sigma$ has Lebesgue measure zero in every chart $\sigma$. In particular, any submanifold of dimension strictly smaller than $N$ has measure zero.
@@ 2.2 Gradient Flows and Asymptotic Behaviour (I)
$df(\mathbf{x})\xi=\langle\operatorname{grad}f(\mathbf{x}),\xi\rangle$ (1)
@@
In local coordinates, the above reads as follows: Let $\sigma:U\to\mathbb{R}^{N}$ be a chart about $\mathbf{x}\in M$, denote $\mathbf{y}=\sigma(\mathbf{x})$, and let $\tilde{f}:=f\circ\sigma^{-1}$ be the chart representation of $f$. Moreover, let $\tilde{\mathsf{g}}:=\sigma_{*}\mathsf{g}$ be the chart representation (where $\sigma_{*}$ denotes the pushforward) of the metric $\mathsf{g}$ on $\sigma(U)$, i.e. $\tilde{\mathsf{g}}(\mathbf{y})=\sum_{i,j=1}^{n}\tilde{\mathsf{g}}_{ij}(\mathbf{y})\,dy^{i}dy^{j}$, where $\tilde{G}(\mathbf{y}):=\big{(}\tilde{\mathsf{g}}_{ij}(\mathbf{y})\big{)}$ is a positive definite matrix varying smoothly in $\mathbf{y}\in\sigma(U)$. Finally, let $\tilde{F}:=\sigma_{*}F$ be the chart representation of the gradient vector field $-\operatorname{grad}f$. Then it holds that
$\tilde{F}^{j}(\mathbf{y})=-\sum_{i=1}^{n}\tilde{\mathsf{g}}^{ij}(\mathbf{y})\frac{\partial\tilde{f}}{\partial y_{i}}(\mathbf{y}),$
where $\tilde{\mathsf{g}}^{ij}(\mathbf{y})$ are the entries of the inverse of $\tilde{G}(\mathbf{y})$ and $\frac{\partial}{\partial y_{i}}$ denotes the $i$-th partial derivative.
Any convergence analysis of gradient systems is based on the following two observations: (i) the critical points of $f$ are the equilibria of (2), (ii) $f$ constitutes a Lyapunov function for (2), i.e. $f$ is monotonically decreasing along solutions. Despite their simplicity, they immediately allow several non-trivial statements about the asymptotic behavior of (2), which can be characterized by its $\omega$-limit sets
@@
If f has compact sublevel sets, i.e. if the sets $\{\mathbf{x}\in M:f(\mathbf{x})\leq c\}$ are compact for all $c\in\mathbb{R}$, then every trajectory of (2) converges to the set of critical points. More precisely, every trajectory of (2) exists for all $t\geq 0$ and its $\omega$-limit set is a non-empty, compact and connected subset of the set of critical points of $f$.
Although Prop. 2.1 shows that $\omega(\mathbf{x}_{0})$ is contained in the set of critical points of $f$, it does not guarantee convergence to a single critical point and, indeed, there are smooth gradient systems whose trajectories exhibit a non-trivial convergence behavior to the set of critical points (cf. “Mexican hat counter-example” by [22]). Yet for isolated critical points, one has the following trivial consequence of Prop. 2.1.
@@
If f has compact sublevel sets and all critical points are isolated, then any solution of (2) converges (for $t\to\infty$) to a single critical point of $f$. Moreover, if additionally all saddle points are strict,555see Subsec. 2.3 below then for almost all initial points (in the sense of Sec. 2.1), the flow converges to a local minimum.